OpenStudy (anonymous):

can someone help me with this problem: fives times a number decreased by 18 minus 4 times the same number is -36.

OpenStudy (anonymous):

5 (x - 18) - 4x = -36 x = -36 + 90 x = 54

OpenStudy (anonymous):

or it might be 5x - 18 - 4x = -36 x = -36 + 18 x = -16

OpenStudy (anonymous):

because im not certain of interpertation of your wording